OpenStudy (anonymous):

17=-5(3+w)+7 answer right and get a medal!

OpenStudy (anonymous):

w=-5

OpenStudy (texaschic101):

17 = -5(3 + w) + 7 --multiply -5 through the parenthesis) 17 = - 15 - 5w + 7 -- get the 5w on the other side, so add 5w to both sides) 17 + 5w = -15 - 5w + 5w + 7 (combine like terms) 17 + 5w = - 8 (now get the 17 to the other side by subtracting 17) 17 - 17 + 5w = -8 - 17 5w = - 25 (divide by 5 to get w by itself) 5w/5 = -25/5 w = - 5